(7x10^4)X/3=4x10^2
Solve for X

Answer by Boreal(15235)

7X10^4* X= 4X 10^2
Divide both sides by 7x10^4
Convert 4x10^2 to 40*10
40*10/7 x 10^4=
5.714 X10^-3
0.00571.
;
Alternative
70000X=400
Divide both sides by 70000
x=400/70,000
x=0.00571